# The search for Population III: Confirmation of a HeII emitter with no metal lines at z=10.6

### 24 Mar 2026 v2

@@ Abstract
We report the confirmation of a HeII1640 emitter located at 3 pkpc from the galaxy GN-z11, at z=10.6. The detection, based on JWST NIRSpec-IFU high-resolution spectroscopy, confirms a previous claim based on medium-resolution spectroscopy. The HeII1640 identification is further supported by the independent detection of H obtained by Übler et al. (2026) at the same location. The HeII emission is spectrally resolved in two components separated by 120 km/s. The Equivalent Width of the HeII emission is extremely high (20Å). No metal lines are detected. Population III stars appear to be the most plausible explanation for the observed HeII emission. We also discuss the possible contribution from a Direct Collapse Black Hole, or a Primordial Black Hole — these scenarios are less plausible, but cannot be ruled out completely.
@@ 7 The Population III scenario
Fig.5 shows the EW(HeII1640) vs HeII/H diagnostic diagram. This diagram was already presented in Maiolino et al. (2024b) for Hebe. In that paper, however, H was measured along the shutter using the prism, hence likely only grazing at the more diffuse emission and not probing the core of the Hebe (see discussion in Section 6). With the new data presented in Übler et al. (2026), we have a direct measurement of H in the same region and with a similar spectral resolution as for the HeII emission. The various symbols show models from Nakajima and Maiolino (2022), with various ionization parameters and metallicities. Given the complete absence of metal lines, to avoid overcrowding the plot, we focus on low metallicity models, below 10% solar. Specifically, red stars indicate PopII models, with decreasing metallicity from dark to light, down to a metallicity of (i.e. ). Note that these models use BPASS binary stellar templates, which are already more energetic than standard stellar populations (see Nakajima and Maiolino, 2022, for details). Orange pentagons are AGN models, with a broad range of possible ionizing spectra, a broad range of ionizing parameters, and a very broad range of metallicities, from 10% solar to completely pristine. The squares show the PopIII models, with sizes indicating the adopted IMF (assuming a Salpeter slope): larger symbols are for stellar masses in the range , intermediate size symbols are for stellar masses in the range , and small symbols for stellar masses in the range . The blue squares are for gas densities of . The purple squares are for gas densities of and . The small blue points are PopIII models from Rusta et al. (2026): dark points are for Pure III pristine systems, while light points are for systems with some PopII contamination, but still dominated by PopIII (50%). The blue points are for gas densities of and ionization parameter , the purple points are for gas densities of and ionization parameter .
@@ 10 The direct collapse black hole or primordial black hole scenarios
Regardless of the potential nature and origin, an accreting black hole in such a metal-poor environment would be an extremely interesting finding. However, there are a few aspects that make these scenarios less plausible. First, the lower limits on the EW(HeII) are uncomfortably high for the AGN scenario, as illustrated in Fig.5; even stronger tension is found by Übler et al. (2026) when exploring the EW(H). Secondly, the observed HeII line profile is quite different from the H one. While BLR stratification can broaden HeII with respect to Balmer lines (e.g. Brazzini et al., 2026), HeII in AGNs is generally not strongly blueshifted, nor it presents a double-peaked profile, unless the Balmer lines also do.
@@ 12 Summary and Conclusions
A Direct Collapse Black Hole or a Primordial Black Hole can potentially be invoked to explain the HeII emission. However, these scenarios face some issues that make them less plausible, although they cannot be entirely ruled out.
@@
In summary, Hebe represents one of the most convincing pieces of evidence for Population III stars in the early Universe, and its properties support models of their formation and early evolution. However, equally compelling alternatives, such as a Direct Collapse Black Hole or a Primordial Black Hole, cannot be completely excluded.
@@ Acknowledgments
We are grateful to Richard Ellis, Harley Katz and Jorryt Matthee for useful comments. This work is based on observations made with the National Aeronautics and Space Administration (NASA)/European Space Agency (ESA)/Canadian Space Agency (CSA) JWST. The data were obtained from the Mikulski Archive for Space Telescopes at the STScI, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programme PID 5086. RM, FD, JS, IJ, GJ acknowledge support from the Science and Technology Facilities Council (STFC), by the European Research Council (ERC) through Advanced Grant 695671 “QUENCH”, by the UK Research and Innovation (UKRI) Frontier Research grant RISEandFALL. RM also acknowledges support from a Royal Society Research Professorship grant. AJB, GC and JC acknowledge funding from the “FirstGalaxies” Advanced Grant from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 789056). HÜ thanks the Max Planck Society for support through the Lise Meitner Excellence Program. HÜ acknowledges funding by the European Union (ERC APEX, 101164796). MP acknowledges support through the grants PID2021-127718NB-I00, PID2024-159902NA-I00, and RYC2023-044853-I, funded by the Spain Ministry of Science and Innovation/State Agency of Research MCIN/AEI/10.13039/501100011033 and El Fondo Social Europeo Plus FSE+. JW gratefully acknowledges support from the Cosmic Dawn Center through the DAWN Fellowship. The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant No. 140. PGP-G acknowledges support from grant PID2022-139567NB-I00 funded by Spanish Ministerio de Ciencia, Innovación y Universidades MCIU/AEI/10.13039/501100011033,FEDER Una manera de hacer Europa. JAAT acknowledges support from the Simons Foundation and JWST program 3215. Support for program 3215 was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. BRP acknowledges support from grants PID2021-127718NB-I00 and PID2024-158856NA-I00 funded by Spanish Ministerio de Ciencia e Innovación y Universidades MCIU/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. DJE acknowledges support by JWST/NIRCam contract to the University of Arizona, NAS5-02105, and as a Simons Foundation Investigator. EE, KH and CNAW acknowledge support by the JWST/NIRCam contract to the University of Arizona NAS5-0215. SC, GV, and SZ acknowledge support by European Union’s HE ERC Starting Grant No. 101040227 - WINGS. WMB gratefully acknowledges support from DARK via the DARK fellowship. This work was supported by a research grant (VIL54489) from VILLUM FONDEN. JMH acknowledges support from JWST Program #8544. Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.

### 27 Mar 2026 v3

@@ Abstract
We report the confirmation of a HeII1640 emitter located at 3 pkpc from the galaxy GN-z11, at z=10.6. The detection, based on JWST NIRSpec-IFU high-resolution spectroscopy, confirms a previous claim based on medium-resolution spectroscopy. The HeII1640 identification is further supported by the independent detection of H obtained by Übler et al. (2026) at the same location. The HeII emission is spectrally resolved in two components separated by 120 km/s. The Equivalent Width of the HeII emission is extremely high (20Å). No metal lines are detected. We argue that Population III stars are the most plausible explanation for the observed He II emission, with no satisfactory alternative from other classes of sources or mechanisms.
@@ 7 The Population III scenario
Fig.5 shows the EW(HeII1640) vs HeII/H diagnostic diagram. This diagram was already presented in Maiolino et al. (2024b) for Hebe. In that paper, however, H was measured along the shutter using the prism, hence likely only grazing at the more diffuse emission and not probing the core of Hebe (see discussion in Section 6). With the new data presented in Übler et al. (2026), we have a direct measurement of H in the same region and with a similar spectral resolution as for the HeII emission. The various symbols show models from Nakajima and Maiolino (2022), with various ionization parameters and metallicities. Given the complete absence of metal lines, to avoid overcrowding the plot, we focus on low metallicity models, below 10% solar. Specifically, red stars indicate PopII models, with decreasing metallicity from dark to light, down to a metallicity of (i.e. ). Note that these models use BPASS binary stellar templates, which are already more energetic than standard stellar populations (see Nakajima and Maiolino, 2022, for details). Orange pentagons are AGN models, with a broad range of possible ionizing spectra, a broad range of ionizing parameters, and a very broad range of metallicities, from 10% solar to completely pristine. The squares show the PopIII models, with sizes indicating the adopted IMF (assuming a Salpeter slope): larger symbols are for stellar masses in the range , intermediate size symbols are for stellar masses in the range , and small symbols for stellar masses in the range . The blue squares are for gas densities of . The purple squares are for gas densities of and . The small blue points are PopIII models from Rusta et al. (2026): dark points are for Pure III pristine systems, while light points are for systems with some PopII contamination, but still dominated by PopIII (50%). The blue points are for gas densities of and ionization parameter , the purple points are for gas densities of and ionization parameter .
@@ 10 The direct collapse black hole or primordial black hole scenarios
Regardless of the potential nature and origin, an accreting black hole in such a metal-poor environment would be an extremely interesting finding. However, there are a few aspects that make these scenarios less plausible. First, the lower limits on the EW(HeII) are uncomfortably high for the AGN scenario, as illustrated in Fig.5. Even stronger tension is found by Übler et al. (2026) when exploring the H emission, as they obtain a lower limit on EW(H) that is an order of magnitude higher than expected by accreting black hole models. Secondly, the observed HeII line profile is quite different from the H one. While BLR stratification can broaden HeII with respect to Balmer lines (e.g. Brazzini et al., 2026), HeII in AGNs is generally not strongly blueshifted, nor it presents a double-peaked profile, unless the Balmer lines also do.
@@ 12 Summary and Conclusions
We have also explored the possibility of a Direct Collapse Black Hole or a Primordial Black Hole. However, these scenarios face significant issues. In particular, they struggle to reproduce the observed high EW(HeII) and EW(H), as well as the emission line profiles.
@@
In summary, Hebe represents one of the most convincing pieces of evidence for Population III stars in the early Universe, and its properties support models of their formation and early evolution.
@@ Acknowledgments
We are grateful to Richard Ellis, Jorryt Matthee, and Daniel Schaerer for useful comments. This work is based on observations made with the National Aeronautics and Space Administration (NASA)/European Space Agency (ESA)/Canadian Space Agency (CSA) JWST. The data were obtained from the Mikulski Archive for Space Telescopes at the STScI, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programme PID 5086. RM, FD, JS, IJ, GJ acknowledge support from the Science and Technology Facilities Council (STFC), by the European Research Council (ERC) through Advanced Grant 695671 “QUENCH”, by the UK Research and Innovation (UKRI) Frontier Research grant RISEandFALL. RM also acknowledges support from a Royal Society Research Professorship grant. AJB, GC and JC acknowledge funding from the “FirstGalaxies” Advanced Grant from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 789056). HÜ thanks the Max Planck Society for support through the Lise Meitner Excellence Program. HÜ acknowledges funding by the European Union (ERC APEX, 101164796). MP acknowledges support through the grants PID2021-127718NB-I00, PID2024-159902NA-I00, and RYC2023-044853-I, funded by the Spain Ministry of Science and Innovation/State Agency of Research MCIN/AEI/10.13039/501100011033 and El Fondo Social Europeo Plus FSE+. JW gratefully acknowledges support from the Cosmic Dawn Center through the DAWN Fellowship. The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant No. 140. PGP-G acknowledges support from grant PID2022-139567NB-I00 funded by Spanish Ministerio de Ciencia, Innovación y Universidades MCIU/AEI/10.13039/501100011033,FEDER Una manera de hacer Europa. JAAT acknowledges support from the Simons Foundation and JWST program 3215. Support for program 3215 was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. BRP acknowledges support from grants PID2021-127718NB-I00 and PID2024-158856NA-I00 funded by Spanish Ministerio de Ciencia e Innovación y Universidades MCIU/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. DJE acknowledges support by JWST/NIRCam contract to the University of Arizona, NAS5-02105, and as a Simons Foundation Investigator. EE, KH and CNAW acknowledge support by the JWST/NIRCam contract to the University of Arizona NAS5-0215. SC, GV, and SZ acknowledge support by European Union’s HE ERC Starting Grant No. 101040227 - WINGS. WMB gratefully acknowledges support from DARK via the DARK fellowship. This work was supported by a research grant (VIL54489) from VILLUM FONDEN. JMH acknowledges support from JWST Program #8544. Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.